OpenStudy (anonymous):

how do you find the real cube root of 216

OpenStudy (amistre64):

i usually just multiply 6*6*6 :)

OpenStudy (amistre64):

gasp...thats satans number.....do another problem lol

OpenStudy (anonymous):

so how would i write the answer 6^3?

OpenStudy (anonymous):

Like this:\[216=6^3\]

OpenStudy (anonymous):

Oh, didn't read it.

OpenStudy (amistre64):

id write it like this: 6 the real cube root of 216 is 6

OpenStudy (anonymous):

\[\sqrt[3]{216}=6\]

OpenStudy (anonymous):

okaie thanks. but how would you find the real cube root of 1000/27?

OpenStudy (amistre64):

calculators the easy way; by hand youd have to do x^3 +3x^2y +3xy^2 +y^3

OpenStudy (amistre64):

1000 = 10^3 27 = 3 ^3 10/3

OpenStudy (anonymous):

ohh. okaiee. Thank yu so much.!

OpenStudy (amistre64):

youre welcome :) just need to memorize those basic cubes.... like the times tables had to be :)

OpenStudy (anonymous):

yeah, that'll take a whilee. lol.